# Euclid Quick Data Release (Q1)

### v1

@@ 2.2.2 Previously known real lenses
Prior to this lens search, there were a number of previously known lenses coinciding with areas of Euclid coverage, which were ideal for training on as they most closely resemble the lenses in Q1. A data set of approximately previously known grade A and B lenses that had Q1 data available was used in training and/or validation of the machine-learning models, but was not counted in the number of newly discovered lenses reported in this paper. This collection of lenses includes lenses discovered in the Euclid ERO data (Acevedo Barroso et al. 2024b), the expert visual inspection of galaxies with spectroscopic velocity dispersion (Euclid Collaboration: Rojas et al. 2025), the Galaxy Zoo Cosmic Dawn project, other surveys that overlapped with Q1, and other serendipitous discoveries. Outside Q1, there were around 50 additional known grade A/B lenses in Euclid discovered from the Galaxy Zoo Euclid project, where volunteers searched the EWS, giving a total number of approximately real Euclid lenses available for training. Around the time of the internal Q1 release, more lenses were discovered serendipitously, which were not discovered in time to add to our training sample, but which are collated in the Q1 lens data set presented in Euclid Collaboration: Walmsley et al. (2025).
As the number of known lenses is small and heterogenous due to the number of different searches they originate from, training on these few known lenses alone may amplify any selection bias in the pre-existing lens sets. Therefore, these lenses were primarily used to supplement the simulated lenses, which are designed to cover a larger range of lensing configurations. Some machine-learning models were trained on the combinations of simulations and real lenses, while others were trained on simulations alone and the real lenses were used to validate their performance. For details of how each machine-learning model was trained see Sect. 4.
@@ 2.2.3 Non-lenses
There were two sets of random cutouts, one of objects in Q1 with the same selection function as the LRGs used for S2 (see Sect. 2.2.1) containing objects, and one of objects with the same selection function as Q1 but from EWS tiles that lie outside the Q1 area. There was a set of labelled non-lenses, which primarily originated from a visual inspection of objects with Dark Energy Spectroscopic Instrument (DESI) data available (Euclid Collaboration: Rojas et al. 2025). This provided a data set of approximately spiral galaxies, mergers, ring galaxies, LRGs, and others. The DESI LRGs were the LRGs used for S1 and S3. Table 1 shows all the data sets available along with which machine-learning models used which sets.
@@ 3.1 Model description
Zoobot222https://github.com/mwalmsley/zoobot (Walmsley et al. 2023) is a Bayesian CNN that is pretrained on 92 million morphological classifications of galaxies as part of the Galaxy Zoo citizen science project (Lintott et al. 2008). Using pretrained models such as Zoobot and fine-tuning them is an example of ‘transfer learning’, where by taking a tool that is already primed to solve a different but related problem and then adapting only a part of the model, it is possible to create a model that is optimised to solve a new problem in a way that is quicker and easier than creating a model from scratch (Walmsley et al. 2024). In our case, we take the pretrained base Zoobot CNN model with pre-determined weights and fine-tune it by freezing all the parameters except those in the final layers. These parameters in the final layers are allowed to vary through training the CNN on new data such as images of strong lenses. The idea of using Zoobot as a base model for lens finding was first presented and tested by Pearce-Casey et al. (2024), where it was found to be the best-performing machine-learning model of the 20 models tested on Euclid ERO data; this motivated the use of Zoobot for lens finding in Q1.
The base Zoobot model was not built to detect lenses originally, nor was it developed on Euclid data. However, lens finding in Euclid is still a morphological classification task using astronomical images. Many of the features of the latent space that Zoobot uses to distinguish the features of galaxy morphology are likely to be relevant to lens finding. While Zoobot will have seen very few images of lenses, it has been pretrained on huge numbers of non-lenses that are likely to be common in the Euclid data set, making it primed to detect these. Pretrained models are especially powerful in tasks such as strong lens finding where the lack of known lenses to train on is a significant issue; fine-tuning a model requires significantly less training data to learn from than a usual model would, and while the number of simulated lenses available is much greater than the number of real lenses, it is still limited. Using a pretrained model also greatly reduces the chance of overfitting as the model is already primed to pick out useful morphological features; Zoobot should already be able to infer what parts of an image are relevant for classification and what to ignore as background objects, artefacts, or stars. By keeping the first layers from the pretrained Zoobot, this also ensures that the model is encouraged to learn true morphological properties rather than other patterns that may arise in simulated lenses such as PSF orientation (Wilde et al. 2022).
@@ 3.2 Fine-tuning approach
There were many different variables that could be adjusted to improve the performance of Zoobotwhen fine-tuning it from a general morphology classifier to a strong lens finder. These included the choice of training data, the choice of what types of images (scalings and filters) to use, and machine-learning hyperparameters (such as batch size and the number of layers fine-tuned). The framework that we used to build an optimal version of Zoobot involved selecting one choice for each of the mentioned options, fine-tuning this version of the model on the choice of training data, and then evaluating each version by its performance on real lenses and non-lenses. This consisted of 97 known grade A Euclid lenses and a data set of non-lens Euclid cutouts from EWS tiles outside Q1 (see Sect. 2.2.3). Since the ratio of 97: does not reflect the proportion of lenses expected in real Euclid data, we scale the statistics to match the ratio of expected lenses according to Collett (2015). For each fine-tuned version of Zoobot, we rank order all of the cutouts. To compare the different versions, we plot (e.g., see Fig. 2) as a function of Zoobot rank the cumulative fraction of the 97 lenses recovered against the total number of images up to that Zoobot rank. This allows us to forecast the number of lenses we should find in Q1 as a function of how many objects we have to inspect.
@@ 3.3 Results from fine-tuning
To quantify the performance of a model, statistics that are of particular interest are the true positive rate (TPR), false positive rate (FPR) and purity (),333TPR is also known as recall or sensitivity; purity is also known as precision. given in terms of true positives (, number of lenses that were predicted to be lenses), false positives (, number of non-lenses that were predicted to be lenses), true negatives (, number of non-lenses that were predicted to be non-lenses), and false negatives (, number of lenses that were predicted to be non-lenses):
@@
TPR can be thought of as the fraction of real lenses that the machine-learning model predicted to be lenses, FPR as the fraction of non-lenses that the machine-learning model predicted to be lenses, and purity as the fraction of things that the machine-learning model predicted to be lenses that are real lenses. Since machine-learning models do not generally produce a binary (lens/non-lens) output, but rather give objects a continuous score from 0 to 1 of how likely an object is to be a lens, the purity, TPR, and FPR vary according to a score threshold at which to split the objects into lenses and non-lenses. There is generally a trade-off between purity and TPR; by setting a high score threshold, fewer objects are classified as lenses and so the proportion of total lenses found is lower, but as the score threshold is lower the purity decreases. The purity at different TPR values for the best performing version of Zoobot is displayed in Fig. 3. As before, these statistics are calculated based on the performance on the known lenses and a set of random Q1 cutouts, with the purity scaled to take into consideration the expected ratio of lenses to non-lenses in such a sample. We recover an almost entirely pure sample of of the lenses, but the purity falls to to recover up to 60% of the known lenses. As can be seen in Fig. 2, the best version of Zoobot requires the first 10% of all the data (score threshold in the range –) to be inspected to reach a TPR of 100%.
@@
When Zoobot was applied to the Euclid ERO data in Pearce-Casey et al. (2024), Zoobot was found to perform better than the other models tested. However, they found that the purity of the lenses discovered was low, which may be partly due to the fact that the number of fine-tuning layers was set to one and was not allowed to vary. Additionally, since the Euclid ERO lens-finding experiment there has been significant development in simulations of Euclid lenses (e.g., Euclid Collaboration: Rojas et al. 2025), which is likely to also contribute to a discrepancy in performance between Euclid ERO and Q1 lens searches. Other factors, such as the limited amount of available Euclid data at the time, training only on random cutouts as negatives instead of a curated set of non-lenses, as well as subsequent improvements to the Zoobot base models (Walmsley et al. 2024), may also contribute to this.
@@
Figure 4 displays random Q1 images at different samples of rankings according to Zoobot. The model is unable to perfectly split the sample into lenses and non-lenses, with non-lenses still contaminating the highly ranked images, meaning that visual inspection is still an essential part of finding strong lenses. However, it is reassuring that at the higher ranks, not only is the density of lenses higher, but it also appears that the lenses that are found are typically more obvious lenses, whether that is due to a high signal-to-noise ratio or a large Einstein radii creating a clear distinction between the lens and arc light. We do not find that the false positives are dominated by spirals, which has historically been a problem for lens finding, and this is likely due to the fact that Zoobot had been pretrained to recognise a diversity of spiral morphologies. Rather, we find that other than a few artefacts, most of the false positives are chance alignment of galaxies; examples of highly ranked false positives by Zoobot are shown in Appendix C.
@@ 4 Other models
In addition to Zoobot, four other models were developed to find lenses in Euclid Q1. In this section we provide a brief overview of each model and refer the reader to the original papers in which they were developed. The models are numbered as follows: Model 1 is a 4-layer CNN; Model 2 is OU-100; Model 3 is IncNet; Model 4 is Zoobot (previously described); and Model 5 is LensVision. This does not provide an exhaustive list of all the machine-learning models that are employed to find lenses in Euclid in general; additional networks were used in the lens search within the ERO data (Pearce-Casey et al. 2024), but were not applied to the Q1 data due to time limitations.
4.1 Model 1: 4-Layer CNN
@@
Model 1 is a shallow 4-layer CNN, with the architecture used being an adaptation of the network developed for the morphological classification of galaxies in Domínguez Sánchez et al. (2018). It has already been tested with Euclid-like simulations (Manjn-Garca 2021) and with the ERO fields (Pearce-Casey et al. 2024; Nagam et al. 2025). The network has four convolutional layers of different spacings (, and , respectively), and two fully-connected layers. ReLU activation functions are applied after every convolutional layer, and a max-pooling is applied after the second and third convolutional layers.
@@
The model was trained on a combination of S1, S2, and S3, and also 80 known grade A and B Euclid lenses as the set of positives. For the negatives, a combination was used of random Q1 galaxies matching the selection function of S2, all the LRGs selected for S1, the labelled DESI non-lenses, and also 7378 cutouts of random Euclid galaxies from non-Q1 tiles. Two versions of the model were trained, both using -only images with fluxes normalised from 0 to 1, but one used the MTF scaling and the other used the scaling, with the final scores being the average of these two outputs.
@@
The training of the network was carried out following a supervised learning approach. Decisions regarding pre-processing, architecture, and learning parameters were influenced by what was found to work when the same model was applied to the ERO data in Pearce-Casey et al. (2024). In the learning process, a binary cross-entropy loss function was used and was optimised by the Adam stochastic optimisation method (Kingma & Ba 2015), with a learning rate of . The models were trained for 40 epochs, using a batch size of 32. In order to prevent over-fitting, several data augmentation techniques were performed during training, allowing the images to be zoomed in and out (0.75 to 1.3 times the original size), rotated, and flipped and shifted both vertically and horizontally. The output of the final convolution layer was flattened and was inputted to a fully-connected layer with dropout probability . The final fully-connected layer used a sigmoid activation to turn the output scores into lens probabilities.
@@ 4.2 Model 2: OU-100
The OU-100 model is another CNN-based model. The network is a modification of OU-200 (Wilde et al. 2022), removing the final two convolutional layers due to the change in input size between the two models ( versus pixels). The architecture consists of four convolutional layers of sizes , and , each followed by a max pooling layer, a dropout layer with dropout probability and a ReLU activation function. This is followed by a fully-connected layer with 350 neurons, followed by an output layer with a softmax activation function to produce a lens prediction.
@@
When inspecting the top 100 ranked candidates from the -only MTF and -only images, it was found that the MTF version appeared better at finding lenses, while lots of the high-ranked objects by the -only version of the model contained stars. In the end the MTF -only version of OU-100 were used.
@@ 4.3 Model 3: IncNet
The IncNet architecture is based on the ‘inception module’ (Szegedy et al. 2015) that applies convolutions of different size on the images in parallel, to extract features on different scales simultaneously and then combines them. The images are passed through two convolutional layers, alternated with max-pooling layers. After this, they go through seven inception modules. The fifth of them is connected to a classifier whose score is combined with the score of the final classifier (i.e., that at the end of the seven inception modules) to guide the weights update (refer to Szegedy et al. 2014 for a more detailed discussion). The output of the classifiers are weighted differently: 0.3 for the intermediate one; and 1.0 for the final one. This baseline architecture, presented in Euclid Collaboration: Leuzzi et al. (2024), was further modified for the Q1 lens finding challenge to account for the input image sizes of pixels compared to the original image sizes it was optimised for, requiring a reduction in the number of filters per layer, while keeping the same number of layers.
For network training, a combination of S1 and S2 was used for the positive class. In the case of S2, the images were augmented by applying either a random rotation or flip to every image. For the negative class, a combination of random galaxies matching S2 and labelled DESI non-lenses were used, with the total number of training images being approximately . For validating the network, a combination of known grade A and B lenses were used, in addition to a non-overlapping set of negatives from the training data. The images used were the -only images with the scaling. During training, an ADAM optimiser was used with an initial learning rate of and binary cross-entropy as the loss function. The model was trained for 100 epochs, and although early-stopping was enabled in the training loop to prevent overfitting, it was not needed.

### v2

@@ 2.2.2 Previously known real lenses
Prior to this lens search, there were a number of previously known lenses coinciding with areas of Euclid coverage, which were ideal for training on as they most closely resemble the lenses in Q1. A data set of approximately previously known grade A and B lenses that had been visually inspected with available Euclid data was used in training and validation of the machine-learning models, but was not counted in the number of newly discovered lenses reported in this paper. This collection of lenses includes lenses discovered in the Euclid ERO data (Acevedo Barroso et al. 2025b), the expert visual inspection of galaxies with spectroscopic velocity dispersion (Euclid Collaboration: Rojas et al. 2025), the Galaxy Zoo Cosmic Dawn project, other surveys that overlapped with Q1, and other serendipitous discoveries. Outside Q1, there were around 50 additional known grade A+B lenses in Euclid discovered from the Galaxy Zoo Euclid project, where volunteers searched the EWS, giving a total number of approximately real Euclid lenses available for training. Around the time of the internal Q1 release, more lenses were discovered serendipitously, which were not discovered in time to add to our training sample, but which are collated in the Q1 lens data set presented in Euclid Collaboration: Walmsley et al. (2025a).
As the number of known lenses is small and heterogeneous due to the number of different searches they originate from, training on these few known lenses alone may amplify any selection bias in the pre-existing lens sets. Therefore, these lenses were primarily used to supplement the simulated lenses, which are designed to cover a larger range of lensing configurations. Some machine-learning models were trained on the combinations of simulations and real lenses, while others were trained on simulations alone and the real lenses were used to validate their performance. For details of how each machine-learning model was trained see Sect. 4.
@@ 2.2.3 Non-lenses
There were two sets of random cutouts, one of objects in Q1 with the same selection function as the deflectors used for S2 (see Sect. 2.2.1) containing objects, and one of objects with the same selection function as Q1 but from EWS tiles that lie outside the Q1 area. There was a set of labelled non-lenses, which originated from a visual inspection of objects with DESI data available (Euclid Collaboration: Rojas et al. 2025). These were a by-product of the visual inspection of high velocity dispersion galaxies that primarily functioned as a lens search, since lensing cross section scales as velocity dispersion to the fourth power, and additionally collected LRGs that could be used for lens simulations. This provided a data set of approximately spiral galaxies, mergers, ring galaxies, LRGs, and others. The DESI LRGs were the LRGs used for S1 and S3. In the future, other catalogues of visually inspected Euclid objects will also be available for machine learning training, such as that from Euclid Collaboration: Walmsley et al. (2025b), or the objects classified as non-lenses from this work. Table 1 shows all the data sets available along with which machine-learning models used which sets.
@@ 3.1 Model description
Zoobot333https://github.com/mwalmsley/zoobot (Walmsley et al. 2023) is a Bayesian CNN that is pretrained on 92 million morphological classifications of galaxies as part of the Galaxy Zoo citizen science project (Lintott et al. 2008). Using pretrained models such as Zoobot and fine-tuning them is an example of ‘transfer learning’, whereby taking a tool that is already primed to solve a different but related problem and then adapting only a part of the model, it is possible to create a model that is optimised to solve a new problem in a way that is quicker and easier than creating a model from scratch (Walmsley et al. 2024). In our case, we take the pretrained base Zoobot CNN model with pre-determined weights and fine-tune it by freezing all the parameters except those in the final layers. These parameters reside in the last layer (the classification head) and the ConvNeXt blocks that precede it, and are allowed to vary through training the CNN on new data such as images of strong lenses. The idea of using Zoobot as a base model for lens finding was first presented and tested by Pearce-Casey et al. (2025), where it was found to be the best-performing machine-learning model of the 20 models tested on Euclid ERO data; this motivated the use of Zoobot for lens finding in Q1.
The base Zoobot model was not built to detect lenses originally, nor was it developed on Euclid data. However, lens finding in Euclid is still a morphological classification task using astronomical images. Many of the features of the latent space that Zoobot uses to distinguish the features of galaxy morphology are likely to be relevant to lens finding. While Zoobot will have seen very few images of lenses, it has been pretrained on a large number of non-lenses that are likely to be common in the Euclid data set, making it primed to detect these. Pretrained models are especially powerful in tasks such as strong lens finding where the lack of known lenses to train on is a significant issue; fine-tuning a model requires significantly less training data to learn from than a usual model would, and while the number of simulated lenses available is much greater than the number of real lenses, it is still limited. Using a pretrained model also greatly reduces the chance of overfitting as the model is already primed to pick out useful morphological features; Zoobot should already be able to infer what parts of an image are relevant for classification and what to ignore as background objects, artefacts, or stars. By keeping the first layers from the pretrained Zoobot, this also ensures that the model is encouraged to learn true morphological properties rather than other patterns that may arise in simulated lenses such as PSF orientation (Wilde et al. 2022).
@@ 3.2 Fine-tuning approach
There were many different variables that could be adjusted to improve the performance of Zoobotwhen fine-tuning it from a general morphology classifier to a strong lens finder. These included the choice of training data, the choice of what types of images (scalings and filters) to use, and machine-learning hyperparameters (such as batch size and the number of layers fine-tuned). The framework that we used to build an optimal version of Zoobot involved selecting one choice for each of the mentioned options, fine-tuning this version of the model on the choice of training data, and then evaluating each version by its performance on real lenses and non-lenses. This consisted of 97 known grade A Euclid lenses and a data set of non-lens Euclid cutouts from EWS tiles outside Q1 (see Sect. 2.2.3). Since the ratio of 97: does not reflect the proportion of lenses expected in real Euclid data, we scale the statistics to match the ratio of expected lenses according to Collett (2015). For each fine-tuned version of Zoobot, we rank order all of the cutouts. To compare the different versions, we plot (e.g. see Fig. 5) the cumulative fraction of the 97 lenses recovered against the total number of images up to that Zoobot rank. This allows us to forecast the number of lenses we should find in Q1 as a function of how many objects we have to inspect.
@@ 3.3 Results from fine-tuning
To quantify the performance of a model, statistics that are of particular interest are the true positive rate (TPR), false positive rate (FPR), and purity (),444The TPR is also known as recall or sensitivity; purity is also known as precision. given in terms of true positives (, number of lenses that were predicted to be lenses), false positives (, number of non-lenses that were predicted to be lenses), true negatives (, number of non-lenses that were predicted to be non-lenses), and false negatives (, number of lenses that were predicted to be non-lenses):
@@
The TPR can be thought of as the fraction of real lenses that the machine-learning model predicted to be lenses, FPR is the fraction of non-lenses that the machine-learning model predicted to be lenses, and purity is the fraction of things that the machine-learning model predicted to be lenses that are real lenses. Since machine-learning models generally do not produce a binary (lens or non-lens) output, but rather give objects a continuous score from 0 to 1 of how likely an object is to be a lens, the purity, TPR, and FPR vary according to a score threshold at which to split the objects into lenses and non-lenses. There is generally a trade-off between purity and TPR; by setting a high score threshold, fewer objects are classified as lenses and so the proportion of total lenses found is lower, but as the score threshold is lower the purity decreases. The purity at different TPR values for the best performing version of Zoobot is displayed in Fig. 6. As before, these statistics are calculated based on the performance on the known lenses and a set of random Q1 cutouts, with the purity scaled to take into consideration the expected ratio of lenses to non-lenses in such a sample. We recovered an almost entirely pure sample of of the lenses, but the purity falls to 20% when recovering up to 60% of the known lenses. As can be seen in Fig. 5, the best version of Zoobot requires the first 10% of all the data (score threshold in the range –) to be inspected to reach a TPR of 100%. Although the model was allowed to train for up to 100 epochs, early stopping was implemented if there was no improvement for ten consecutive epochs. The best performing version reached an optimal performance after 13 epochs, with each version taking under an hour to train on an NVIDIA A100 80GB GPU.
@@
When Zoobot was applied to the Euclid ERO data in Pearce-Casey et al. (2025), Zoobot was found to perform better than the other models tested. However, they found that the purity of the lenses discovered was low, which may be partly due to the fact that the number of fine-tuning layers was set to one and was not allowed to vary. Additionally, since the Euclid ERO lens-finding experiment there has been significant development in simulations of Euclid lenses (e.g. Euclid Collaboration: Rojas et al. 2025), which is likely to also contribute to a discrepancy in performance between Euclid ERO and Q1 lens searches. Other factors, such as the limited amount of available Euclid data at the time, training only on random cutouts as negatives instead of a curated set of non-lenses, as well as subsequent improvements to the Zoobot base models (Walmsley et al. 2024), may also contribute to this.
@@
Figure 7 displays random Q1 images at different samples of rankings according to Zoobot. The model is unable to perfectly split the sample into lenses and non-lenses, with non-lenses still contaminating the highly ranked images, meaning that visual inspection is still an essential part of finding strong lenses. However, it is reassuring that at the higher ranks, not only is the density of lenses higher, but it also appears that the lenses that are found are typically more obvious lenses, whether that is due to a high signal-to-noise ratio or a large Einstein radii creating a clear distinction between the lens and arc light. We do not find that the false positives are dominated by spirals, which has historically been a problem for lens finding. This is likely due to the fact that Zoobot had been pretrained to recognise a diversity of spiral morphologies, along with the fine-tuning on high-resolution space-based imaging. Rather, we find that other than a few artefacts, most of the false positives are chance alignment of galaxies; examples of highly ranked false positives by Zoobot are shown in Appendix C.
@@ 4 Other models
In addition to Zoobot, four other models were developed to find lenses in Euclid Q1. In this section we provide a brief overview of each model and refer the reader to the original papers in which they were developed. The models are numbered as follows: Model 1 is a four-layer CNN; Model 2 is OU-100; Model 3 is IncNet; Model 4 is Zoobot (previously described); and Model 5 is LensVision. There were no performance criteria for a model to be used for Q1 since it was hard to extrapolate from any pre-Q1 data how well these models would perform in practice. Rather, these models represent all the models that had been trained and were available at the time. This does not provide an exhaustive list of all the machine-learning models that are employed to find lenses in Euclid in general; additional networks were used in the lens search within the ERO data (Pearce-Casey et al. 2025), but were not applied to the Q1 data due to time limitations.
4.1 Model 1: Four-layer CNN
@@
Model 1 is a shallow four-layer CNN, with the architecture used being an adaptation of the network developed for the morphological classification of galaxies in Domínguez Sánchez et al. (2018). It has already been tested with Euclid-like simulations (Manjn-Garca 2021) and with the ERO fields (Pearce-Casey et al. 2025; Nagam et al. 2025). The network has four convolutional layers of different spacings (, and , respectively), and two fully connected layers. ReLU activation functions are applied after every convolutional layer, and a max-pooling is applied after the second and third convolutional layers.
@@
The model was trained on a combination of S1, S2, and S3, and also 80 known grade A and B Euclid lenses as the set of positives. For the negatives, we used a combination of random Q1 galaxies matching the selection function of S2, all the LRGs selected for S1, the labelled DESI non-lenses, and also 7378 cutouts of random Euclid galaxies from non-Q1 tiles. Two versions of the model were trained, both using -only images with fluxes normalised from 0 to 1, but one used the MTF scaling and the other used the scaling, with the final scores being the average of these two outputs.
@@
The training of the network was carried out following a supervised learning approach. Decisions regarding pre-processing, architecture, and learning parameters were influenced by what was found to work when the same model was applied to the ERO data in Pearce-Casey et al. (2025). In the learning process, a binary cross-entropy loss function was used and was optimised by the Adam stochastic optimisation method (Kingma & Ba 2015), with a learning rate of . The models were trained for 40 epochs, using a batch size of 32. In order to prevent overfitting, several real-time data augmentation techniques were performed randomly at each epoch during training. These augmentations included allowing the images to be zoomed in and out (0.75 to 1.3 times the original size), rotated, and flipped and shifted both vertically and horizontally. The output of the final convolution layer was flattened and was inputted to a fully connected layer with dropout probability . The final fully connected layer used a sigmoid activation to turn the output scores into lens probabilities.
@@ 4.2 Model 2: OU-100
The OU-100 model is another CNN-based model. The network is a modification of OU-200 (Wilde et al. 2022), removing the final two convolutional layers due to the change in input size between the two models ( versus pixels). The architecture consists of four convolutional layers with kernels of sizes , and , each followed by a max pooling layer, a dropout layer with dropout probability , and a ReLU activation function. This is followed by a fully connected layer with 350 neurons, followed by an output layer with a softmax activation function to produce a lens prediction.
@@
The top 100 ranked images of the four versions of the model were visually inspected. For the + MTF and + MTF versions, the model picked out a few stars, as well as other features such as diffraction spikes and artefacts. The rest of the images were spirals or lens candidates. When comparing the top 100 ranked candidates from the -only MTF and -only images, it was found that the MTF version appeared better at finding lenses. The -only version had 34 stars in the top 100 images, as well as many very clear spirals. Based on the number of lens candidates and false positives in the top 100 images of each version, the MTF -only version of the model performed the best, and this was used as the final model.
@@ 4.3 Model 3: IncNet
The IncNet architecture is based on the ‘inception module’ (Szegedy et al. 2016) that applies convolutions of different size on the images in parallel, to extract features on different scales simultaneously and then combines them. The images are passed through two convolutional layers, alternated with max-pooling layers. After this, they go through seven inception modules. The fifth of them is connected to a classifier whose score is combined with the score of the final classifier (i.e. that at the end of the seven inception modules) to guide the weights update (refer to Szegedy et al. 2015 for a more detailed discussion). The output of the classifiers are weighted differently: 0.3 for the intermediate one; and 1.0 for the final one. This baseline architecture, presented in Euclid Collaboration: Leuzzi et al. (2024), was further modified for the Q1 lens finding challenge to account for the input image sizes of pixels compared to the original image sizes it was optimised for, requiring a reduction in the number of filters per layer, while keeping the same number of layers.
For network training, a combination of S1 and S2 was used for the positive class. In the case of S2, the images were augmented by applying either a random rotation or flip to every image. For the negative class, a combination of random galaxies matching S2 and labelled DESI non-lenses were used, with the total number of training images being approximately . For validating the network, a combination of known grade A and B lenses were used, in addition to a non-overlapping set of negatives from the training data. The images used were the -only images with the scaling. During training, an ADAM optimiser was used with an initial learning rate of , the ReduceLROnPlateau method was implemented to reduce the learning rate in case of overfitting, and binary cross-entropy was used as the loss function. The model was trained for 100 epochs, and although early stopping was enabled in the training loop to prevent overfitting, it was not needed.